OpenStudy (anonymous):

Um, im having serious problems with this question and I would gladly appreciate it if one of you could EXPLAIN to me how to solve it, it is 37= -4s - 1.

OpenStudy (shadowfiend):

To solve this, you want to get s on its own.

OpenStudy (anonymous):

ok, I'm sorry.

OpenStudy (shadowfiend):

So we'll rearrange it so that the s is on the left: -4s - 1 = 37

OpenStudy (shadowfiend):

Then we think about how to get rid of that one. If we add 1 to both sides, the left side ones will turn into 0: -4s - 1 + 1 = 37 + 1 -4s - 0 = 38 -4s = 38

OpenStudy (shadowfiend):

Now we see that if we divide both sides by -4, we end up with just an s on the right, because the -4s cancel on the right: \[\begin{align} \frac{-4s}{-4} &= \frac{38}{-4}\\ \frac{\cancel{-4}s}{\cancel{-4}} &= -\frac{38}{4}\\ s &= -\frac{38}{4}\\ \end{align}\] We can simplify this fraction by noticing that 4 is 2 * 2: \[\begin{align} s &= -\frac{38}{2\cdot 2}\\ s &= -\frac{\cancel{38} 19}{\cancel{2} \cdot 2}\\ s &= -\frac{19}{2} \end{align}\] We cancel the 2 out with half of the 38, leaving us with 19/2.

OpenStudy (anonymous):

Thank you very much, sir! :3